OpenStudy (anonymous):

Q2. You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your head from the top of the tree in a straight line?

OpenStudy (onepieceftw):

It's a right triangle. Use the Pythagorean theorem. a^2 + b^2 =c^2 so take 120^2 + 50^2 = x and solve for x. After you find x take it's root and that's your answer.